A college sold a total of 1000 tickets for a concert.Lower level tickets were $9 and upper level tickets were $5.If they took in $6200,how many of each type of ticket did they sell?5
Answer by checkley71(8403)

L+U=1000 OR L=100-U
9(1000-U)+5U=6200
9000-9U+5U=6200
-4U=6200-9000
-4U=-2800
U=-2800/-4
U=700 UPPER LEVEL TICKETS SOLD.
L+700=1000
L=1000-700
L=300 LOWER TICKETS SOLD.
PROOF:
9(300)+5*700-6200
2700+3500=6200
6200=6200